Opened the roof of the observatory around 10pm; telescope and the mount will cool down to the ambient temperature now.
An AAVSO alert just came in: A request for multiband photometry of a star in Taurus: "V0725 Tau".
All the equipment got connected easily, except the Pyxis 3 Rotator.. tried everything, unplugging and plugging back, reinstalling the drivers, restarting the Sky X Pro software and re starting windows. Nothing worked! Finally found out, the port has mysteriously been changed. It is working fine now. It just wasted my one and half an hour.
Camera cooling point is changed tonight. So far i had adjusted it at -20C.. from tonight, it is at -30C!!
An hour was left before the star will cross meridian line.
Found the star chart from AAVSO.. I do have at least one reference star in my field of view.
AAVSO has requested for photometric V filter data as a priority but since moon is so close to the star, i have decided to use photometric I filter, which will cut down a lot of sky background light.
Slewed the telescope to a nearby 5 magnitude star, focused the telescope and returned to my target area.
A guide star of of magnitude ~10.5 is available. After calibrating the autoguider setup with the Pyxis 3 Rotator's Position Angle of 144, the guide star is now being tracked at 0.5 Hz.
Made the folders for the images, adjusted the camera settings with the exposure of 180 seconds and hit the 'Take Series' button in Sky X Pro.
A few images came bad.. it seems the fog has started to build up in the CCD chamber because of very low temperature settings. I would have to bake the camera desiccant very soon.
On 6th image the CCD chip is now doing fine... Imaging, continues..
Ah.. unfortunately.. me being careless, did not test the exposure time of the single images. 180 seconds turned out to be too bright for the target star and i somewhat saturated it. In simple words, all the data so far taken is useless now.
I have changed the temperature of the CCD now.. so i need to take dark frames. Doing that now. 30 Seconds darks are done.. 60 seconds in progress.
30, 60 and 90 seconds dark frames have been recorded. Roof closed now!